OpenStudy (anonymous):

Given two points on the line, input the equation of the line in standard form Ax + By = C. Reduce all fractional answers to lowest terms. (7, -3), (4, -8)

OpenStudy (anonymous):

no problem any doubt in any more question

OpenStudy (anonymous):

Input the equation for the following line in standard form Ax + By = C. Reduce all fractional answers to lowest terms. A line that passes through the origin and is parallel to x + y = 6.

OpenStudy (anonymous):

find the slope: (-3-8)/(7-4)=5/3 Then plug the slope and a point into the slope intercept equation: -3=(5/3)(7)+b and you get y=(5/3)x-(44/3) Multiply both sides by 3 and then put x and y on the same side to get to standard form and get: 5x-3y=44

OpenStudy (somethingawesome):

For your second question, you can use that same method. We need a point( x,y) and the slope of the line. The point is the origin, so (0,0). The slope is going to be the same as the slope in the line x+y=6, since that's what it means to be parallel to that line.

OpenStudy (anonymous):

Input the equation for the following line in standard form Ax + By = C. Reduce all fractional answers to lowest terms. A line with y-intercept of 2 that is parallel to 2x + y = -5.

OpenStudy (anonymous):

subtract 2x so you have y=2x-5 which is in slope intercept form: y=mx+b m is the slope and b is the y intercept. Since it is parallel you know the slope is the same and it tells you the y intercept is 2 so you keep the slope the same and change be from -5 to 2 and get y=2x+2

OpenStudy (anonymous):

Input the equation for the following line. Reduce all fractional answers to lowest terms. A line that passes midway between and is parallel to both x - y = 6 and x - y = -6.

OpenStudy (aroub):

y=-x

OpenStudy (anonymous):

Given two points on the line, input the equation of the line in standard form Ax + By = C. Reduce all fractional answers to lowest terms. (-8, 0), (1, 5)

OpenStudy (aroub):

slope is 5-0/1+8 = 5/9 , y=mx+b (1,5) 5=5/9 (1)+b , -5/9 +45/9 =b , 40/9 =b y=5/9x+40/9

OpenStudy (aroub):

oh and for u to learn and pass u try it and we will check it for u (two)!! :D

OpenStudy (anonymous):

Which of the following points lies on the line whose equation is y = 2x - 3?

OpenStudy (aroub):

ok ill help u (two)

OpenStudy (aroub):

to find the slope u will have like this (x1,y1) (x2,y2) <-- the slope format will become y2-y1/x2-x1

OpenStudy (aroub):

^ did u understand???

OpenStudy (anonymous):

no thats not a choice on the test lol

OpenStudy (anonymous):

Which of the following points lies on the line whose equation is y = 2x - 3? (0, 3) (1, -1) (-1, 4)

OpenStudy (aroub):

ashley and tiffany im trying to help u !!! ssheh

OpenStudy (anonymous):

we dont know how to do it lol just tell us the answer lmao